OpenStudy (anonymous):

let A and B be events such that Pr (A U B)= .9 and Pr [A ]=.5. What is Pr [B] in the following case that A and B are independent?

OpenStudy (anonymous):

if they are independent than Pr (A U B)= Pr [A +Pr [B] hope ts helps

OpenStudy (anonymous):

this is wrong sorry, got mixed up had a feeling that this is the case

OpenStudy (anonymous):

P(A) + P(B) - P(AnB) = P(AuB) plus if they are independent than P(AnB) = P(A)P(B) .

OpenStudy (anonymous):

so in this case 0.9=0.5+P(B)-P(AnB)=0.5+P(B)-0.5P(B)

OpenStudy (anonymous):

solving this gives P(B)=0.8

OpenStudy (anonymous):

thanks also another question.. does [(E U F)'] = E' U F' or does it intersect

OpenStudy (anonymous):

For the first one you might find it easier to do: \[P (A \cup B) = P(A) - P\left(\begin{matrix}B \\A\end{matrix}\right)\] which gives \[0.9 = 0.5 + \left(\begin{matrix}B \\0.5\end{matrix}\right)\]

OpenStudy (anonymous):

(EUF)' = E' (intersection) F'